OpenStudy (anonymous):

Solve |2y - 2| = y. {2/3, 2} {2, 3} {2, 5} {1/3, 2}

OpenStudy (anonymous):

I got {2,2/3} but that is not one of the answers

OpenStudy (dtan5457):

It's {2/3, 2}. Want the explanation?

OpenStudy (anonymous):

Yes Please

OpenStudy (ksaimouli):

yup 2/3,2 is right

OpenStudy (dtan5457):

An absolute value equation like this can go 2 ways. One way is: 2y-2=y, you solve for y which means -y=-2, or y=2. That's the first one. For the next one, you negate the equation. -(2y-2)=y -2y+2=y, 3y=2, y= 2/3

OpenStudy (anonymous):

Thanks